HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

ANALYSIS OF THE INFRARED SPECTRUM OF ISOTOPICALLY CHIRAL TRANS-2,3-DIDEUTEROOXIRANE

S. ALBERT, K. KEPPLER, G. WICHMANN, M. QUACK, Department of Chemistry and Applied Biosciences, ETH Zurich, 8093 Zurich, Switzerland; Z. CHEN, College of Chemistry and Chemical Engineering, Lanzhou University,

730000 Lanzhou China; J. STOHNER, ICBC, ZHAW, 8820 Wädenswil Switzerland; V. SCHURIG, Institut für Organische Chemie, Universität Tübingen, 72074 Tübingen, Germany; O. TRAPP, Department of Chemistry, LMU, 80539

München, Germany

The concept of isotopic chirality is of fundamental interest because it introduces a completely new isotope effect arising from the parity violating weak nuclear force^{1}^{,}^{2}^{,}^{3}. It also has possible astrophysical applications. Oxirane (c-C_{2}H_{4}O), detected by astrophysical spectroscopy^{4}, has two isotopically chiral isotopomers, c-C_{2}H_{3}DO and trans-2,3-dideuterooxirane (c-CHD-CHDO). We have previously reported the analysis of high resolution GHz and THz spectra of c-C_{2}H_{3}DO in the context of a possible first astrophysical observation of an isotopically chiral species^{5}^{,}^{6}. Following this work, Müller et al. have reported^{7} a tentative detection of c-C_{2}H_{3}DO and the achiral c-CD_{2}CH_{2}O. There have also been ab initio calculations^{8} on c-C_{2}H_{3}DO and c-C_{2}H_{4}O. We have previously reported^{9} high resolution GHz and THz spectra of trans-c-CHD-CHDO. Here we report results of the analysis of the infrared spectrum of trans-2,3-dideuterooxirane, recorded with a resolution of 0.0015 cm^{−1} using the Bruker IFS 125 HR Zürich Prototype (ZP 2001) Fourier Transform spectrometer. We have used the ground state molecular parameters obtained from our analysis^{9} and have carried out the rovibrational analysis using Watson’s A reduced effective Hamiltonian. The results will be discussed as they pertain to isotopic chirality and molecular parity violation.

1M. Quack, Angew. Chem. Int. Ed. 1989, 28, 571.

2R. Berger, G. Laubender, M. Quack, A. Sieben, J. Stohner, M. Willeke, Angew. Chem. Int. Ed.

2005, 44, 3623.

3M. Hippler, E. Miloglyadov, M. Quack and G. Seyfang, in Handbook of High-resolution Spectroscopy Vol. 2, p. 1069, M. Quack and F. Merkt, Eds., Wiley, Chicester, 2011. See also the recent review by M. Quack, G. Seyfang, G. Wichmann, Chem. Sci. 2022, 13, 10598.

4J. E. Dickens, W. M. Irvine, M. Ohishi, M. Ikeda, S. Ishikawa, A. Nummelin, A. Hjalmarson, ApJ 1997, 489, 753.

5S.Albert,Z.Chen,K.Keppler,P.Lerch,M.Quack,V.Schurig,O.Trapp, PCCP 2019,21,3669.

6S.Albert,Z.Chen,K.Keppler,M.Quack,J.Stohner,V.Schurig,O.Trapp, 2023 [in prep.].

7H. S. P. Müller, J. K. Jørgensen, J.-C. Guillemin, F. Lewen, S. Schlemmer, J. Mol. Spectrosc., 2023, 394, 111777; Monthly Notices of the Royal Astronomical Society 2023, 518(1), 185.

8C. Puzzarini, M. Biczysko, J. Bloino, and V. Barone, ApJ 2014, 785, 107.

9Z. Chen, S. Albert, K. Keppler, M. Quack, V. Schurig, O. Trapp, ISMS 2021, WJ01.

p-number: p222  Submitted on Wed, 02 Aug 2023 11:17:38 +0200